Assume that y varies inversely as x. If y =4 when x = 1/2, find y when x =2
Answer by fractalier(2101)

If y varies inversely as x, we have (as before)
y = k/x
4 = k/(1/2)
k = 2
Now
y = 2/x
y = 2/2
y = 1